OpenStudy (anonymous):

WILL FAN/ MEDAL Simplify. –7 + 8 + 10 ÷ (–2) I think it's -4

OpenStudy (gabebae):

Do you know what PEMDAS is?

OpenStudy (anonymous):

no

OpenStudy (gabebae):

\[-7 + 8 -5\] -4 is your answer

OpenStudy (anonymous):

it is -4,

OpenStudy (gabebae):

PEMDAS is Parentheses, Exponents, Multiplication, Division, Addition, and Subtraction

OpenStudy (gabebae):

It's the order how you solve these types of question

OpenStudy (anonymous):

can ether of you help with other questions? and i do know the order of operations.

OpenStudy (anonymous):

@GabeBae @AskMeLotss

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Simplify. [6.6 ÷ (–5 + 3)] • (–1)

OpenStudy (anonymous):

the answer is 3.3

OpenStudy (anonymous):

what you do is -5+3=-2 6.6/-2=-3.3x-1=3.3 a negative times a negative is always positive

OpenStudy (anonymous):

thanks. Which property could be used to find the missing number? (–8 + 3.1) • (4.2 + 7) = (4.2 + 7) • (–8 + ) A.commutative property <-- ? B.distributive property C.associative property

OpenStudy (anonymous):

just a sec. please

OpenStudy (anonymous):

It's A, because the Commutative property is the condition that a group of quantities connected by operators gives the same result whatever the order of the quantities involved, e.g., a × b = b × a.

OpenStudy (anonymous):

can I get a fan and medal :D

OpenStudy (anonymous):

thanks, :) anymore questions

OpenStudy (anonymous):

100% :D

OpenStudy (anonymous):

:D YAY!

OpenStudy (anonymous):

btw i like your profile picture its cool, i love monkeys

OpenStudy (gabebae):

oh I came late sorry about that